My homework is about making algebraic statements. Can you please help me with "A rectangle is 5 times as long as it is wide. If it is w feet wide, how long is it? What is the perimeter?": My homework is about making algebraic statements. Can you please help me with "A rectangle is 5 times as long as it is wide. If it is w feet wide, how long is it? What is the perimeter?"
Answer by checkley77(3654)

W=WIDTH.
THEN LENGTH=5*W.
PERIMETER=2*L+2*W
PERIMETER=2*5W+2W
PERIMETER=10W+2W
PERIMETER=12W